# A Survey on 3D Egocentric Human Pose Estimation

### v1

@@ 3.2 Model Based Methods
In this section, we expand on the different model based egocentric pose estimation methods, overview of which is provided in Table 3.
Dittadi et al. [58] introduced a method using variational autoencoders to generate human body poses from noisy head and hand pose data. It addresses the challenge of predicting full body poses with limited information by training specialized inference models. The approach involves pre-training a regular VAE with full body poses, followed by optimizing the model using incomplete information. Despite limitations in walking motion representation and assumptions about hand signals, the method proves effective in reconstructing plausible body poses.
Figure 4: An example of human model generation based egocentric pose estimation framework [64]. The model processes egocentric video to predict head pose using a hybrid approach merging monocular SLAM, GravityNet, and HeadNet. Predicted head pose guides a conditional diffusion model for full-body pose generation.
Yuan et al. [57] proposed a novel approach employing control-based modeling with physics simulation and utilizes imitation learning to acquire a video-conditioned control policy for ego-pose estimation. Traditional computer vision methods focus solely on motion kinematics [74] neglecting the underlying physics of dynamics [75]. Taking this into account, this framework allows domain adaptation, transferring the policy from simulation to real-world data.
CoolMoves [59], a pioneering Virtual Reality (VR) system, has achieved real-time, expressive full-body motion synthesis for a user’s avatar using limited input cues from VR systems. It delineates the movements more prominent through dynamic fusion with stylized examples from skilled performers. The system excels in synthesizing upper and lower-body motions without explicit tracking cues, addressing challenges in activities like dancing and fighting.
To solve the problem of top-down view distortions, EgoRenderer [60] offered a novel method tackling the complexities of rendering full-body avatars from egocentric images by decomposing the process into texture synthesis, pose construction, and neural image translation [76, 77]. They introduced a neural network for dense correspondence and texture extraction, and an implicit texture stack for capturing dynamic appearance variations. The approach involves estimating body pose and synthesizing a free-viewpoint pose image, and using a neural image translation network to combine it with textures.
The Human POSEitioning System (HPS) [61] combines wearable sensors, IMUs, and a head-mounted camera to accurately track and integrate 3D human poses within pre-scanned environments. By fusing camera-based localization with IMU-based tracking and scene constraints, HPS achieves drift-free, physically plausible motion estimation.
To address challenges like body truncation and pose ambiguities, [65] proposed a human mesh recovery model. It introduces a scene-conditioned diffusion model guided by a physics-based collision score, facilitating the generation of realistic human-scene interactions. It utilizes classifier-free training for flexibility in sampling, providing accurate estimations for visible body parts and diverse results for unseen parts.
Jiang et al. [62] developed a pioneering learning-based approach that predicts full-body poses in world coordinates solely from motion input derived from the user’s head and hands. Leveraging a transformer encoder, the method extracts deep features, distinguishing global motion from local joint orientations to facilitate pose estimation.
FLAG [63] generates realistic 3D human body poses in mixed reality applications using sparse input signals from head mounted devices. It employs a flow-based generative model to predict full-body poses and provides uncertainty estimates for joints. It enhances prior work by generating high-quality poses from limited data and enabling efficient sampling of latent variables for pose generation and optimization.
Su et al. [68] proposed a novel method to estimate 3D wearer poses from egocentric video, overcoming challenges of body invisibility and complex motion. The approach involves a data processing framework converting raw video to 3D pose, incorporating a lightweight Self-Perception Excitation (SPE) module for self-understanding from egocentric view.
EgoEgo [64] was introduces for 3D human motion estimation from monocular egocentric videos. The method, as shown in Figure 4, estimates ego-head pose and generates ego-body pose, allowing independent learning without paired datasets. It combines monocular SLAM and transformer-based models for accurate ego-head pose estimation, employing a conditional diffusion model for full-body pose generation based on the predicted head pose. The innovation lies in leveraging head motion as an intermediate representation, enabling the use of diverse datasets effectively.
Cuevas-Velasquez et al. proposed SimpleEgo [69] which is a novel approach for regression of probabilistic full-body pose parameters from head-mounted camera images. It directly predicts joint rotations, eliminating the need for iterative fitting processes or manual tuning. By representing joint rotations as matrix Fisher distributions, the model predicts confidence scores, allowing for robust uncertainty estimation. The contributions include probabilistic joint rotation prediction, in-depth uncertainty analysis, and a prospect of real-world deployment.
EgoPoser [66], a novel approach introduced for full-body pose estimation using sparse motion sensors, focusing on HMD-based ego-body pose estimation in large scenes. It addresses overfitting issues in existing methods by emphasizing position-invariant prediction with a Global-in-Local motion decomposition strategy. Notably, EgoPoser adapts to diverse body sizes, remains robust when hands are out of view, and reduces motion artifacts.
4 Evaluation Metrics
Several approaches have been introduced to assess 3D egocentric human pose estimation. In this section, we delve into the evaluation metrics commonly used to assess egocentric pose estimation across different skeletal and model based methods.
MPJPE (Mean Per Joint Position Error): This metric is widely utilized to measure the accuracy of predicted 3D joint positions in comparison to ground truth positions. The calculation of MPJPE involves determining the Euclidean distance between the estimated 3D joints and their corresponding ground truth positions. Mathematically, it is expressed as follows:
$\text{MPJPE}=\frac{1}{N}\sum_{i=1}^{N}\|J_{i}-J^{*}_{i}\|$ (1)
Here, $N$ is the total number of joints, $J_{i}$ denotes the estimated position of the $i$th joint, $J_{i}^{*}$ represents the ground truth position of the $i$th joint. MPJPE serves as a robust measure to quantify the overall accuracy of 3D pose estimation algorithms across different joints.
@@ 4 Evaluation Metrics
Table 4: Comparison of different 3D egocentric pose estimation methods on Mo2Cap2 dataset using MPJPE (mm).
@@
Table 5: Comparison of different 3D egocentric pose estimation methods on xr-EgoPose dataset using MPJPE (mm).
@@
BA-MPJPE involves resizing bone lengths to a standard skeleton and then calculating Procrustes-aligned MPJPE between the sequences. This metric provides a comprehensive evaluation by considering structural consistency in bone lengths and eliminating body scale influence.
Global MPJPE aligns all poses within a batch to ground truth, considering translation and rotation and evaluates global joint position accuracy.
MPJRE (Mean Per Joint Rotation Error) calculates the average rotational disparity between predicted and ground truth joint rotations.
MPJVE (Mean Per Joint Velocity Error) assesses the average velocity difference between predicted and ground truth joint movements.
Percentage of Correct Key-points (PCK) metric is a measure of accuracy that checks if the predicted keypoint and the actual joint are close within a specific distance limit. Typically, this distance threshold is set based on the size of the subject.
Head Orientation Error assesses rotational accuracy in head pose estimation. It measures the angular disparity by calculating the Frobenius norm of the difference between the predicted and ground truth head rotation matrices.
Head Translation Error focuses on translation accuracy. It quantifies the mean Euclidean distance between predicted and ground truth head trajectories, providing a quantitative measure of spatial disparity.
@@ 5 Performance Analysis
Refer to captionInput ImageMo2Cap2Mo2Cap2 + Wang et al. [20]xr-EgoPosexr-EgoPose + Wang et al. [20]Hwang et al. [45]
@@
Figure 5: Qualitative comparison between different state-of-the-art egocentric 3D pose estimation models on xr-EgoPose dataset [24]. The predicted 3D poses (highlighted in red) are superimposed onto the ground truth poses (depicted in blue).
This section aims to provide a comprehensive evaluation of various egocentric pose estimation methods across multiple datasets. This in-depth analysis is crucial for understanding the versatility and adaptability of these methods in different scenarios. Each dataset may present unique challenges such as varied lighting conditions, backgrounds, and subject movements. We analyze the performance of different state-of-the-art methods for the 3D egocentric human pose estimation on some of these popular egocentric datasets.
5.1 Performance on Mo2Cap2 dataset
Mo2Cap2 [22] dataset has been widely used in egocentric pose estimation approaches for training and evaluation purpose using the MPJPE metric. The performance of different egocentric pose estimation methods across different actions on this dataset have been depicted in Table 4. The average MPJPE across all actions reduces from 97.85 mm in VNect [78] method to 46.61 mm in SelfPose-UNet [6] method. We see that 2D-3D lifting models [6, 24] achieved better results than direct 3D pose estimation methods. It may be due to the preserved uncertainty information of the joints. The methods evaluated on this dataset do not perform well on the poses where body parts are partially visible or occluded. The Mo2Cap2 [22] is a synthetic dataset having a large number of training and testing data. Though the dataset comprises 3000 subject actions and 700 body textures, performance of these methods deteriorates significantly in complex real world scenarios.
5.2 Performance on xr-EgoPose dataset
The xr-EgoPose [24] dataset has a large number of dataset comprising 383k frames but the action class is quite low compared to the Mo2Cap2 [22] dataset. It has been extensively used by many state-of-the-art egocentric pose estimation methods for evaluating their models. From Table 5, we see that EgoGlass [25] outperforms all the existing methods in terms of mean joint error. It may be Sdue to utilizing body part information for low visible joints. Figure 5 demonstrates a visual representation of 3D pose generated by different state-of-the-art egocentric 3D pose estimation models from same egocentric inputs from xr-EgoPose dataset [24]. The mean joint error is less than the methods on Mo2Cap2 [22] dataset for actions with less visible joints. But, in this dataset, most of the actions that are used for evaluation are quite simple and conspicuous. Due to this, these actions may not be a proper benchmark to evaluate the 3D pose estimation from egocentric view.
5.3 Performance on other datasets
Apart from the mentioned datasets, there also exists some noteworthy datasets which have been used for evaluation in egocentric pose estimation. Table 6 and Table 7 depicts the results of different methods in UnrealEgo [17] and Wang et al.’s [20] dataset demonstrating MPJPE and PA-MPJPE. We can see that methods on UnrealEgo [17] dataset perform comparatively better than the methods evaluated on Wang et al.’s [20] dataset. It is due to 17 different action in diversified environments with a large amount of indoor and outdoor scenes in UnrealEgo [17] dataset whereas Wang et al.’s [20] dataset contains few classified actions and scenes. From the results in the mentioned tables, we observe that xr-EgoPose [24] performs worse in both the datasets. This maybe due to the evaluating datasets being real world whereas xr-EgoPose [24] is trained in synthetic dataset. Camera setup can also be an issue in this case because this method fails significantly in the out-of-field view.
Moreover, datasets like ECHA, EgoGlass, EgoCap etc. have also been used by various egocentric pose estimation models for evaluation and ablation study purpose. In summary, the performance evaluation section provides valuable insights into the efficacy of various egocentric pose estimation methods across diverse datasets.
Methods MPJPE PA-MPJPE
xR-egopose [24] 112.86 88.71
EgoGlass [25] 86.45 63.71
UnrealEgo [17] 79.06 59.95
Ego3DPose [50] 60.82 48.47
Akada et al. [30] 50.55 40.50
EgoTap [55] 41.06 35.39
@@ 5.3 Performance on other datasets
Table 6: Comparison of 3D Egocentric Pose Estimation methods on UnrealEgo [17] dataset.
Methods MPJPE PA-MPJPE
xR-egopose [24] 112.0 87.20
Mo2Cap2 [22] 102.3 74.46
Rhodin et al. [81] 89.67 73.56
EgoPW [18] 81.71 64.87
Scene Aware [19] 76.50 61.92
Table 7: Comparison of 3D Egocentric Pose Estimation methods on Wang et al.’s [20] dataset.
6 Discussion and Future Directions
In this survey paper, we provide a systematic overview of 3D egocentric human pose estimation, encompassing diverse datasets and estimation methodologies. We highlight recent studies addressing the challenge of estimating 3D human pose from an egocentric viewpoint using RGB images or video sequences, an area that remains relatively under-explored in research. Consequently, researchers have recently proposed diverse datasets with lightweight setups. However, the absence of standardized benchmark datasets poses a challenge for evaluating the robustness of different egocentric pose estimation models. Recent advancements, especially by the introduction of Ego-Exo4D [29], offers promise as a valuable resource to address this research gap.
@@ 6 Discussion and Future Directions
We provide a systematic overview of several skeletal and model based methods for egocentric pose estimation, along with a discussion of their individual strengths and weaknesses. We realize that most of the existing methods encounter difficulties for pose estimation in case of in-the-wild scenarios. This is primarily due to insufficient in-the-wild training data which affects the generalization ability of different egocentric pose estimation models. Notably, similar to traditional pose estimation, the biggest challenges of egocentric pose estimation models are strong occlusions and limited field of view, especially for the lower body joints. Multi view consistency may help to to solve this using additional 3D information. Moreover, temporal and contextual information can be utilized further to improve the robustness of the models considering these issues. Consequently, there exists ample scope for refining egocentric pose estimation approaches to better suit real-time XR-technologies.
@@
In conclusion, this survey paper serves as a comprehensive resource for researchers seeking to explore the existing egocentric pose estimation methods, understand prevalent challenges, and make further advancements in this important field.
Acknowledgements

### v2

@@ 3.2 Body Shape Based Methods
In this section, we expand on the different human body shape-based egocentric pose estimation methods found in the literature. Out of them, 12 are highlighted in Table 3.
@@
Table 4: Comparison of different skeletal based 3D egocentric pose estimation methods on Mo2Cap2 dataset using MPJPE ($m$).
@@
Table 5: Comparison of different skeletal based 3D egocentric pose estimation methods on xR-EgoPose dataset using MPJPE ($m$).
Dittadi et al. [14] used variational autoencoders to generate human body poses from noisy head and hand pose data. It addresses the challenge of predicting full body poses with limited information by training specialized inference models. Yuan et al. [71] employed control-based modeling with physics simulation and used imitation learning to acquire a video-conditioned control policy for ego-pose estimation. Traditional computer vision methods focus solely on motion kinematics [42] neglecting the underlying physics of dynamics [43]. Taking this into account, this framework allows domain adaptation, transferring the policy from simulation to real-world data. CoolMoves [2] is a VR system that has achieved real-time, expressive full-body motion synthesis for a user’s avatar using limited input cues from VR systems. It delineates the prominent movements through dynamic fusion with stylized examples from skilled performers. The system excels in synthesizing upper and lower-body motions without explicit tracking cues, addressing challenges in activities like dancing and fighting.
To solve the problem of top-down view distortions, EgoRenderer [21] renderes full-body avatars from egocentric images by decomposing the process into texture synthesis, pose construction, and neural image translation [52, 58]. The Human POSEitioning System (HPS) [20] combines wearable sensors, IMUs, and a head-mounted camera to accurately track and integrate 3D human poses within pre-scanned environments. By fusing camera-based localization with IMU-based tracking and scene constraints, HPS achieves physically plausible motion estimation.
To address challenges like body truncation and pose ambiguities, [75] introduced a scene-conditioned diffusion model guided by a physics-based collision score, facilitating the generation of realistic human-scene interactions. It uses classifier-free training for flexibility in sampling, providing accurate estimations for visible body parts and diverse plausible results for unseen parts. [30] predicted full-body poses in world coordinates solely from motion input derived from the user’s head and hands. Leveraging a transformer encoder, the method extracts deep features, distinguishing global motion from local joint orientations to facilitate pose estimation. FLAG [5] uses sparse input signals from head mounted devices and a flow-based generative model to predict full-body poses and provide uncertainty estimates for joints. [54] estimated 3D wearer poses from egocentric video, overcoming challenges of body invisibility and complex motion. They convert raw video to 3D pose, incorporating Self-Perception Excitation module for self understanding from egocentric view.
EgoEgo [35] uses monocular egocentric videos to estimate ego-head pose and generate ego-body pose, allowing independent learning without paired datasets. It combines monocular SLAM and transformer-based models for accurate ego-head pose estimation, employing a conditional diffusion model for full-body pose generation based on the predicted head pose. SimpleEgo [11] performs regression of probabilistic full-body pose parameters from head-mounted camera images. It directly predicts joint rotations, eliminating the need for iterative fitting processes or manual tuning. By representing joint rotations as matrix Fisher distributions, the model predicts confidence scores, allowing for robust uncertainty estimation. AGRoL [15] proposes a lightweight MLP-based diffusion model for realistic full-body motion synthesis from sparse tracking signals. [34] introduces affordable motion capture using smartwatches and head-mounted camera, integrating head poses for sparsity, tracking floor levels for outdoor settings, and optimizing motion with visual cues. EgoPoser [31] generates full-body pose estimation using sparse motion sensors, focusing on HMD-based ego-body pose estimation in large scenes. It addresses overfitting issues by emphasizing position-invariant prediction with a Global-in-Local motion decomposition strategy. Notably, it adapts to diverse body sizes and remains robust when hands are out of view.
4 Evaluation Metrics
In this section, we briefly describe the different metrics used to assess 3D egocentric human pose estimation methods.
MPJPE (Mean Per Joint Position Error) is a widely utilized metric which measures the mean error between all the predicted 3D joint positions and the ground truth positions, by calculating the Euclidean distance between them.
@@ 4 Evaluation Metrics
BA-MPJPE first resizes the bone lengths to a standard skeleton and then calculates the PA-MPJPE, providing a comprehensive evaluation by considering structural consistency in bone lengths and eliminating body scale influence.
Global MPJPE evaluates global joint position accuracy by aligning all poses within a batch to the ground truth, considering translation and rotation.
MPJRE (Mean Per Joint Rotation Error) and MPJVE (Mean Per Joint Velocity Error) compares the predicted and ground truth joints by calculating the average rotational and velocity disparity respectively.
Percentage of Correct Key-points (PCK) is a measure of accuracy that checks if the predicted keypoint and the actual joint are close within a specific distance limit. Typically, this distance threshold is set based on the size of the subject.
Head Translation & Orientation Error focuses on translation and rotational accuracy in head pose estimation respectively. The translation error is quantified using the mean Euclidean distance between predicted and ground truth head trajectories. Whereas, the orientation error is calculated using the Frobenius norm of the difference between the predicted and ground truth head rotation matrices.
@@ 5 Performance Analysis
In this section, we compare the performance of different state-of-the-art methods for the 3D egocentric human pose estimation on some of the popular egocentric datasets.
Performance of Skeletal-based Methods: Table 4 shows the performance of five different skeletal-based egocentric pose estimation methods across the different actions on the widely used Mo2Cap2 [68] dataset. The average MPJPE across all actions reduces from 66.8 mm in EgoFish3D [39] method to 46.61 mm in SelfPose-UNet [61] method. We see that 2D-3D lifting models [61, 60] achieved better results than direct 3D pose estimation methods, which may be due to the preserved uncertainty information of the joints. Figure 3 shows the qualitative evaluation of the 3D poses generated by three different skeletal-based methods on the xR-EgoPose [60] dataset. Table 5 compares six different skeletal based methods on the xR-EgoPose [60] dataset. Overall, the MPJPE of the methods is lower here than those tested in Mo2Cap2 [68] dataset, especially for actions with less visible joints. This could be because in this dataset most of the actions used for evaluation are relatively simpler.
Refer to captionInput ImageMo2Cap2 [68]xR-EgoPose [60]Hwang et al. [25]
@@
Figure 3: Qualitative comparison between three different state-of-the-art skeletal-based egocentric 3D pose estimation models on the xR-EgoPose dataset [60]. The predicted 3D poses (red) are superimposed onto the ground truth poses (blue).
Methods MPJPE MPJVE
CoolMoves [2] 7.83 100.54
Lee et al. [34] 5.87 19.11
AvatarPoser [30] 4.18 29.40
EgoPoser [31] 4.14 25.95
AGRoL [15] 3.86 50.94
AGRoL-Offline [15] 3.71 18.59
@@
Table 6: Comparison of different body shape based 3D Egocentric Pose Estimation methods on AMASS [44] dataset using MPJPE (cm) and MPJVE (cm/s).
Performance of body shape based Methods: Table 6 shows the performance of six different body shape based egocentric pose estimation methods on the AMASS [44] dataset. We can see that, AGRoL [15] outperforms other methods with its smooth motion generation, but it’s limited to offline use. For real-time applications, EgoPoser [31] is more suitable as it provides more adaptability to diverse body sizes as well as robustness with hands out of view. Figure 4 qualitatively compares the 3D human pose and shape on three different body shape based methods using HPS [20] dataset.
Refer to captionGroundtruth ImageAvatarPoser [30]AgRoL [15]EgoPoser [31]
Figure 4: Qualitative comparison of three different state-of-the-art body shape based egocentric 3D pose estimation models on the HPS dataset. [20].
6 Conclusion and Future Directions
@@
In this survey paper, we provide an overview of 3D egocentric human pose estimation using RGB images or video sequences, encompassing diverse datasets and estimation methodologies. Researchers have proposed diverse datasets with lightweight setups. However, the lack of standardized benchmark datasets, except for the recent Ego-Exo4D [18] dataset, poses a challenge for evaluating the robustness of different egocentric pose estimation models. While discussing the individual strengths and weaknesses of different skeletal and body shape based methods for egocentric pose estimation, we realize that most of the existing methods encounter difficulties with in-the-wild scenarios mainly due to insufficient training data. Notably, similar to traditional pose estimation, the biggest challenges of egocentric pose estimation models are strong occlusions and limited field of view, especially for the lower body joints. Multi view consistency may help to to solve this using additional 3D information. Moreover, temporal and contextual information can be utilized further to improve the robustness of the models considering these issues. Consequently, there exists ample scope for refining egocentric pose estimation approaches to better suit real-time technologies.
@@
In conclusion, this survey paper serves as a comprehensive resource for researchers seeking to explore the existing egocentric pose estimation methods, understand prevalent challenges, and make further advancements.
7 Acknowledgements